In order to determine the effects of smoking on lung cancer, 250 individuals were surveyed, including 75 who had lung cancer. Of those who smoked, 70 had lung cancer and of those without cancer, 165 did not smoke.

a) Declare the categorical variables and their categories. Make a contingency table.
b) What percentage of individuals smoke and do not have lung cancer?
c) What percentage of individuals do not smoke?
d) What percentage of individuals who have lung cancer do not smoke?
Answer #1

Ans:

a)

Categorical variables are:

smoke or don't smoke

lung cancer or No lung cancer

lung cancer No lung cancer Total
smoke 70 10 80
don't smoke 5 165 170
Total 75 175 250

b)

Percentage of individuals smoke and do not have lung cancer=100*10/250=4%

c)

Percentage of individuals do not smoke=100*170/250=68%

d)

Percentage of individuals who have lung cancer do not smoke=100*5/75=6.67%